# references.bib

### v5

@@ 5 Covariance between order flow imbalance and prices changes
An interesting quantity that can be computed within our model and easily measured empirically using the public tape of trades and prices is the “aggregated” impact $\mathbb{E}[\Delta|I^{a}]$, conditioned to a certain value of imbalance $I^{a}$, as studied in [plerou2002quantifying, patzelt2018universal]. We know that such a quantity behaves very differently from the square-root law, and has non-trivial scaling properties as a function of $T$, see [patzelt2018universal] for $a=0$ and $a=1$. For $a=0$, in particular, one finds that the initial slope of $\mathbb{E}[\Delta|I^{0}]$ as a function of $I^{0}$ scales like $T^{-\omega}$ with $\omega\approx 1/4$ [patzelt2018universal, bouchaud2018trades], a result we confirm in Fig. 7.
@@
5.1 Without volume fluctuations
Since impact fluctuations, considered in section 4.5, are independent of the order imbalance (i.e. $\mathbb{E}[\varepsilon\eta]=0$), we can restrict the calculation to the deterministic part of impact, which is given by the generalized propagator model above. Neglecting any sign correlations between different metaorders, we thus get, when all child orders have the same size:
$\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]=\nu\varphi q^{a}\mathcal{I}_{1}(q,\varphi)\int_{0}^{T}{\rm d}u\int_{0}^{\infty}{\rm d}s\Psi(s)\left[\mathbb{I}(s>u)u\sqrt{u}+\mathbb{I}(s<u)s\sqrt{s}\left((\frac{u}{s})^{1-\beta}-(\frac{u}{s}-1)^{1-\beta}\right)\right].$ (75)
This quantity scales like $T^{1-\beta}$ whenever $\mu>3/2+\beta$ and like $T^{5/2-\mu}$ otherwise. [^11] Note that for the standard propagator model, one finds that the scaling is always $T^{3-\beta-\mu}$ when $\mu>1$. Note that if one disregards the non-Gaussian nature of $I^{0}$ and uses Eq. (74) to compute the initial slope of the aggregate impact, one finds (using Eq. (17)) $\omega=1/2$ when $\mu=3/2$ and $\beta>0$, which is far from the empirical value $\omega\approx 1/4$. As we shall see in the following, empirical results suggest a strong dependence on $a$, meaning that volume fluctuations certainly cannot be neglected in this calculation.
5.2 With correlated metaorders
One can extend the result above to account for correlated metaorders. Assuming $\gamma>0$ and $\beta<1/2$, the “non-diagonal” contribution $t\neq t^{\prime}$ gives a contribution that scales as
$\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]_{o.d.}\propto\Gamma\,q^{a+\frac{1}{2}}\,T^{2-\beta-\gamma_{\times}},$ (76)
to be compared with the above results, i.e. $T^{1-\beta}$ or $T^{5/2-\mu}$. When $\gamma_{\times}<1$, the non-diagonal contriburion is always dominant at large $T$ compared to $T^{1-\beta}$ and becomes subdominant compared to $T^{5/2-\mu}=T^{3/2-\gamma}$ when $\beta+\gamma_{\times}>1/2+\gamma$.
For $\gamma=\gamma_{\times}$ and $\beta<\frac{1}{2}$ we thus deduce that the “off-diagonal” contribution is always dominant for large enough times. The corresponding value of the slope exponent $\omega$ is now equal to $\beta+\gamma_{\times}-\gamma$, and is thus very close to the empirical value $\omega\approx 1/4$ when $\gamma_{\times}\approx\gamma$ and $\beta\approx 0.2$.
5.3 With correlated metaorders and volume fluctuations
Let us now add volume fluctuations, with a $q$-dependent value of $\mu_{q}$ as given by Eq. (22) and $\beta_{q}=\beta_{1}-\lambda^{\prime}\ell$, with $\ell=\log q$. Now, there exists a value $q=q_{c}^{\prime}$ such that $5/2-\mu_{q_{c}^{\prime}}=1-\beta_{q_{c}^{\prime}}$.
Then, using $\mathcal{I}_{1}(q,\varphi)\propto\sqrt{q}$, one gets, for the diagonal contribution
@@ 5.3 With correlated metaorders and volume fluctuations
In section 5.7 below, we will show that empirical data can be fitted as a power-law of $T$, with an effective exponent that indeed behaves non-monotonically with $a$, which suggests $(\lambda,\lambda^{\prime})\sigma_{\ell}^{2}$ in the range 0.10.10.10.10.20.20.20.2, also in line with the condition already obtained in section 3.2.
@@
depending on whether the diagonal or off-diagonal contribution dominates. Numerically, with $\lambda\sigma_{\ell}^{2}=1/8$ and $\gamma_{\times}+\widehat{\beta}(0)=0.8$, one finds $\omega_{d.}\approx 0.56$ and $\omega_{n.d.}\approx 0.30$, close to the empirical value $1/4$ in the second case.
5.4 With a random impact component
Quite interestingly, if the random component of impact (see Eq. (68)) is assumed to be such that $\mathbb{E}[\varepsilon\eta]=0$, it will not contribute to the covariance between price changes and order imbalance. Indeed, by definition such a contribution to price changes does not contribute to the covariance between price changes and order imbalance. In such a scenario, only the fundamental component can contribute to the covariance, provided some metaorders are “informed”, as we show next.
5.5 With “informed” metaorders
Up to now, we assumed that there are no correlations between the sign of metaorders $\varepsilon$ and the “Fundamental” component of price changes on time scale $T$, $\sigma_{F}\xi\sqrt{T}$, see Eq. (68). If we rather assume that $\mathbb{E}[\varepsilon\xi]=\rho/\sqrt{\nu T}$, where $\rho$ measures the average amount of information of individual metaorders, [^12] For a detailed discussion of the scaling with $T$, see [bouchaud2010impact], section 16.1.3. The idea is that out of $N_{T}=\nu T$ metaorders of random signs, an excess fraction $\sim\sqrt{\nu T}$ is possibly informed. It is important to stress that, in the spirit of the Kyle model [kyle1985continuous], the fundamental component $\sigma_{F}\xi\sqrt{T}$ is not a mechanical consequence of impact. we find an extra contribution to $\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]$ computed above, which reads:
@@ 5.5 With “informed” metaorders
5.6 The correlation coefficient
Finally, we turn our attention with a natural description of the interplay between (generalized) volume imbalance and price changes, namely the following correlation coefficient
@@ 5.6 The correlation coefficient
For $a\to 0$, the power of $T$ is very close to zero for our default choice of parameters. For higher values of $a$, the exponent becomes positive and therefore this off-diagonal contribution adds an increasing function of $T$ [^13] In the asymptotic $T\to\infty$ limit, one should take into account the fact that $\Sigma_{a}$ itself becomes dominated by the off-diagonal contribution, see the discussion around Eq. (39). Hence the correlation does saturate when $T\to\infty$, as it should be.
@@
The scaling with $T$ indicates that $R_{a}(T)$ should decrease with $T$ for small values of $a$, and become independent of $T$ for large values of $a$. Note that if volatility is mostly due to fundamentals and not due to impact, one finds that $R_{a}(T)$ is directly proportional to $\rho$ [^14] It may be realistic to assume that well informed metaorders are larger, and therefore that $\rho\propto q^{\psi}$ where $\psi\geq 0$. In this case, the scaling with $a$ reads $R_{a}(T)_{F}\propto e^{{\sigma_{\ell}^{2}}a(2\psi-a)/2}$, which reaches a maximum for $a^{\star}=\psi$.
@@
5.7 Empirical data
@@
    Provided the main source of price moves is the average impact of random metaorders, the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ behaves as a power-law of $T$ with an effective exponent that is non-monotonic in $a$, reaching a minimum for some value of $a$, see Eq. (80). If instead volatility is dominated by the random component of impact (Eq. (68)) with a “Fundamental” component $\sigma_{F}\xi$, a linear behaviour in $T$ independent of $a$ should be observed;
@@
5.7.1 Power-law behaviour of the covariance
[…]
The value of the effective exponent is in the range 0.950.950.950.951.351.351.351.35, as expected since the diagonal contributions give an exponent slightly below $1$ (Eq. (80)) and the off-diagonal contribution yield an exponent $\approx 1.2$ for the default values of $\gamma_{\times}$ and $\widehat{\beta}(0)$ (see Eq. (81)).
@@ 5.7.1 Power-law behaviour of the covariance
5.7.2 Correlation vs. $T$ and $a$
Turning to point 2., Fig. 9 shows the correlation coefficient $R_{a}(T)$ vs. $T$ for different $a$ in two different representations: standard plot and heatmap. A first immediate observation is that these correlations are $O(1)$ for all $T$ values, and peak around $0.45$ for stocks and $0.7$ for futures. This means that order flow and returns are indeed strongly correlated, as was emphasized many times (see e.g. [evans2002order, chordia2004order, hopman2007supply, bouchaud2009markets, patzelt2018universal]).
@@
Finally, note that the saturation regime where $R_{a}(T)$ should become independent of $T$, as predicted by either the “diagonal” hypothesis ((85)) or the “Fundamental” hypothesis (Eq. (87)), is hardly observable in the data, at least up to $10^{4}$ trades (roughly one trading day). This observation appears to confirm the predominant role of impact of mostly uninformed (but correlated) metaorders in the genesis of volatility [bouchaud2018trades].
@@
5.7.4 Empirical covariance: conclusion
All empirical data appear to confirm the qualitative validity of our predictions, some of them being rather non-trivial. One of the main assumptions of our model is that the exponents describing the autocorrelation of child orders ($\mu_{q}$) and the decay of their impact ($\beta_{q}$) depend on the size $q$ of these child orders. This, in turn, leads to a non-monotonic dependence of the scaling of the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ with $a$, [^15] Such a non-monotonic behaviour cannot be explained simply from the power-tail distribution of volumes $q$, see G. Maitrier et al., in preparation. and of the correlation $R_{a}(T)$ with both $T$ and $a$. Although some of our predictions fail to explain the data quantitatively, we are tempted to ascribe these discrepancies on the bluntness of our approximations, which, we argue, correctly capture the mechanisms at play.
@@ 5.7.4 Empirical covariance: conclusion
  • The hypothesis according to which most of the volatility comes from fundamental information is hard to reconcile with the data. First, as discussed in section 4.5, this assumption does not enable one to rationalize why the square-root law, which applies to all metaorders (informed or not [gomes2015market, bucci2018slow, sato2024does]) is proportional to volatility. Second, it predicts that the covariance of price returns and volume imbalances is proportional to $T$ independently of $a$, at odds with empirical data. As argued in [bouchaud2006random, taranto2018linear, bouchaud2018trades], the most plausible hypothesis is that volatility stems from trading alone. In other words, the excess volatility puzzle has a microstructural origin.
@@ 6 Conclusion
In order to answer these questions, we have introduced a new theoretical framework to describe metaorders with different signs, sizes and durations, possibly correlated between themselves, which all impact prices as a square-root of volume (which we assume as an input) but with a subsequent time decay characterized by an exponent $\beta\neq\frac{1}{2}$, i.e. different from the one suggested by the classic propagator model [bouchaud2003fluctuations, bouchaud2018trades] or the LLOB model [donier2015fullyconsistentminimalmodel]. We proposed a generalized propagator model to account for such a feature.
@@
We then established that the power-law tailed distribution of metaorder durations is not sufficient to counteract impact decay that leads to price sub-diffusion. Rather, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders, which is indeed present in data. In fact, we conjecture that the intra- and cross-correlations between child orders decay roughly in the same manner, a feature that may be crucial for explaining the success of the construction of synthetic metaorders from public data [maitrier2025generatingrealisticmetaorderspublic].
@@
The existence of correlations between metaorders is therefore a crucial ingredient to recover price diffusion. The old debate between order splitting and “herding” that seemed to have been closed by several papers in favor of splitting [toth2015equity, sato2023inferring], is perhaps not so clear-cut. Such correlations could be due either to the fact that many participants use the same trading signals, or that copy-cat metaorders/unwinding market maker inventories correlate to past order flow, or else that some traders successfully predict the future behaviour of other participants. Note that within our story any predictive alpha signal manifests itself through autocorrelated metaorders, since only trades can move the price (see also the discussion in [bouchaud2018trades], chapter 20).
@@
In view of the strongly fluctuating volumes $q$ of child orders, one quickly realizes that one needs to account for heterogeneity in the distribution of metaorder durations, and in the resulting decay of their impact, which we parametrized by two $q$-dependent exponents, $\mu_{q}$ and $\beta_{q}$, assumed to depend linearly on $\ell=\log q$. In a nutshell, this is needed to account for the fact that metaorders with large $q$ are less autocorrelated than those with small $q$, as shown in Fig. 3. This feature allowed us to account semi-quantitatively for the way the moments of generalized volume imbalance scale with time $T$, and more importantly how the correlation between price changes and generalized volume imbalance scales with $T$. We predicted that the corresponding power-law should depend in a non-monotonic fashion on the parameter $a$ that allows one to put the same weight on all child orders ($a=0$) or overweight large orders ($a$ large), a behaviour clearly borne out by empirical data, see Fig. 8. We also predicted that the correlation between price changes and volume imbalances should display a maximum as a function of $a$ for fixed $T$, which again matches observations, see Fig. 10, with fitting parameters fixed with previously determined values. We found that stocks and futures appear to differ quite markedly in terms of these metrics, which could provide an interesting new way to characterize price formation mechanisms in different markets.
@@
Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We claim that our results strongly support the “Order-Driven” theory of markets, according to which it is the mechanical impact of trades, independently of any notion of “fundamental information”, that generates volatility in financial markets, a picture advocated in [bouchaud2006random, hopman2007supply, bouchaud2018trades, bouchaud2022inelastic] and, in a different context, in [gabaix2021search]. In particular, the Efficient Market Hypothesis, which posits that volatility is mostly due to the variation of “fundamental value”, cannot easily explain our results concerning the correlations between price changes and order flow imbalance.
@@
All these ideas will be quantitatively confirmed in a follow-up paper (G. Maitrier et al., in preparation) using a numerical simulation of our model. Such a numerical implementation will allow us in to (i) generate realistic synthetic datasets that capture the subtle interplay between order flows and prices, and (ii) validate on such synthetic data the method of metaorder reconstruction from the public tape recently proposed in [maitrier2025generatingrealisticmetaorderspublic].

### v6

@@ 6 Covariance between order flow imbalance and prices changes
An interesting quantity that can be computed within our model and easily measured empirically using the public tape of trades and prices is the “aggregated” impact $\mathbb{E}[\Delta|I^{a}]$, conditioned to a certain value of imbalance $I^{a}$, as studied in [43, 42]. We know that such a quantity behaves very differently from the square-root law, and has non-trivial scaling properties as a function of $T$, see [42] for $a=0$ and $a=1$. For $a=0$, in particular, one finds that the initial slope of $\mathbb{E}[\Delta|I^{0}]$ as a function of $I^{0}$ scales like $T^{-\omega}$ with $\omega\approx 1/4$ [42, 5], a result we confirm in Fig. 6.
@@
6.1 Correlated metaorders and volume fluctuations
In this subsection, we compute the covariance $\mathbb{E}[\Delta\cdot I^{a}]$ in the full blown model, with correlated metaorders and volume fluctuations. As above, we assume a $q$-dependent value of $\mu_{q}$ as given by Eq. (21) and further postulate a similar behaviour for $\beta_{q}$, i.e.
$\beta_{q}=\beta_{1}-\lambda^{\prime}\ell,$ (62)
with $\ell=\log q$ (see also Appendix A).
Now, there exists a value $q=q_{c}^{\prime}$ such that $5/2-\mu_{q_{c}^{\prime}}=1-\beta_{q_{c}^{\prime}}$. Then, using $\mathcal{I}_{1}(q,\varphi)\propto\sqrt{q}$, one gets, for the diagonal contribution
@@ 6.1 Correlated metaorders and volume fluctuations
In section 7 below, we will show that empirical data can be fitted as a power-law of $T$, with an effective exponent that indeed behaves non-monotonically with $a$, which suggests $(\lambda,\lambda^{\prime})\sigma_{\ell}^{2}$ in the range 0.10.10.20.2, also in line with the condition already obtained in section 4.2.
@@
depending on whether the diagonal or off-diagonal contribution dominates. Numerically, with $\lambda\sigma_{\ell}^{2}=1/8$ and $\gamma_{\times}+\widehat{\beta}(0)=0.8$, one finds $\omega_{d.}\approx 0.56$ and $\omega_{o.d.}\approx 0.30$, close to the empirical value $1/4$ in the second case.
6.2 The role of “informed” metaorders
Up to now, we assumed that there are no correlations between the sign of metaorders $\varepsilon$ and the “fundamental” component of price changes on time scale $T$, $\sigma_{F}\xi\sqrt{T}$, see Eq. (A7). If we rather assume that $\mathbb{E}[\varepsilon\xi]=\rho/\sqrt{\nu T}$, where $\rho$ measures the average amount of information of individual metaorders, [^8] For a detailed discussion of the scaling with $T$, see [9], section 16.1.3. The idea is that out of $N_{T}=\nu T$ metaorders of random signs, an excess fraction $\sim\sqrt{\nu T}$ is possibly informed. It is important to stress that, in the spirit of the Kyle model [32], the fundamental component $\sigma_{F}\xi\sqrt{T}$ is not a mechanical consequence of impact. we find an extra contribution to $\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]$ computed above, which reads:
@@ 6.2 The role of “informed” metaorders
6.3 From covariances to correlations
Finally, we turn our attention to a natural description of the interplay between (generalized) volume imbalance and price changes, namely the correlation coefficient
@@ 6.3 From covariances to correlations
For $a\to 0$, the power of $T$ is very close to zero for our default choice of parameters. For higher values of $a$, the exponent becomes positive and therefore this off-diagonal contribution adds an increasing function of TT. [^9] In the asymptotic $T\to\infty$ limit, one should take into account the fact that $\Sigma_{a}$ itself becomes dominated by the off-diagonal contribution, see the discussion around Eq. (38). Hence the correlation does saturate when $T\to\infty$, as it should be.
@@
The scaling with $T$ indicates that $R_{a}(T)$ should decrease with $T$ for small values of $a$, and become independent of $T$ for large values of $a$. Note that if volatility is mostly due to fundamentals and not due to impact, one finds that $R_{a}(T)$ is directly proportional to $\rho$. [^10] It may be realistic to assume that well informed metaorders are larger, and therefore that $\rho\propto q^{\psi}$ where $\psi\geq 0$. In this case, the scaling with $a$ reads $R_{a}(T)_{F}\propto e^{{\sigma_{\ell}^{2}}a(2\psi-a)/2}$, which reaches a maximum for $a^{\star}=\psi$.
@@
7 Covariances & Correlations: Empirical data
@@
    Provided the main source of price moves is the average impact of random metaorders, the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ behaves as a power-law of $T$ with an effective exponent that is non-monotonic in $a$, reaching a minimum for some value of $a$, see Eq. (66). If instead volatility is dominated by the random component of impact (Eq. (A7)) with a “fundamental” component $\sigma_{F}\xi$, a linear behaviour in $T$ independent of $a$ should be observed;
@@
7.1 Power-law behaviour of the covariance
We first investigate point 1. above. As shown in Fig. 7 (left), a power-law behaviour of $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ as a function of $T$ is approximately verified for all $a$, albeit with some amount of convexity (for TSCO) or concavity (for LLOY or EUROSTOXX). When plotted as a function of $a$, the effective exponent of $T$ displays the predicted non-monotonic behaviour, see Fig. 7 (right), reaching a minimum for $a\approx 1$ for LLOY and TSCO, $a\approx 1.1$ for SPMINI and $a\approx 1.5$ for EUROSTOXX. The left slope is predicted to be equal to $\lambda\sigma^{2}_{\ell}$ and the right slope to $\lambda^{\prime}\sigma^{2}_{\ell}$. For LLOY, TSCO and SPMINI we thus find $\lambda\sigma^{2}_{\ell}\approx 0.1$ (not far from the estimate derived from Fig. 3) and $\lambda^{\prime}\sigma^{2}_{\ell}\approx 0.15$$0.2$. For EUROSTOXX, we estimate $\lambda\sigma^{2}_{\ell}\approx 0.5$, a factor 2 larger than from the behaviour of $\Sigma_{a}$ (Fig. 3), but a very small (but positive) value for $\lambda^{\prime}\sigma^{2}_{\ell}$.
The value of the effective exponent is in the range 0.950.951.351.35, as expected since the diagonal contributions give an exponent slightly below 11 (Eq. (66)) and the off-diagonal contribution yield an exponent $\approx 1.2$ for the default values of $\gamma_{\times}$ and $\widehat{\beta}(0)$ (see Eq. (67)).
@@ 7.1 Power-law behaviour of the covariance
7.2 Correlation vs. $T$ and $a$
Turning to point 2., Fig. 8 shows the correlation coefficient $R_{a}(T)$ vs. $T$ for different $a$ in two different representations: standard plot and heatmap. A first immediate observation is that these correlations are $O(1)$ for all $T$ values, and peak around $0.45$ for stocks and $0.7$ for futures. This means that order flow and returns are indeed strongly correlated, as was emphasized many times (see e.g. [21, 16, 29, 6, 42]).
@@
Finally, note that the saturation regime where $R_{a}(T)$ should become independent of $T$, as predicted by either the “diagonal” hypothesis ((71)) or the “fundamental” hypothesis (Eq. (73)), is hardly observable in the data, at least up to $10^{4}$ trades (roughly one trading day). This observation appears to confirm the predominant role of impact of mostly uninformed (but correlated) metaorders in the genesis of volatility [5].
@@
7.4 Empirical covariances & correlations: summary
All empirical data appear to confirm the qualitative validity of our predictions, some of them being rather non-trivial. One of the main assumptions of our model is that the exponents describing the autocorrelation of child orders ($\mu_{q}$) and the decay of their impact ($\beta_{q}$) depend on the size $q$ of these child orders. This, in turn, leads to a non-monotonic dependence of the scaling of the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ with aa, [^11] Such a non-monotonic behaviour cannot be explained simply from the power-tail distribution of volumes $q$, see G. Maitrier et al., in preparation. and of the correlation $R_{a}(T)$ with both $T$ and $a$. Although some of our predictions fail to explain the data quantitatively, we are tempted to ascribe these discrepancies on the bluntness of our approximations, which, we argue, correctly capture the mechanisms at play.
@@ 7.4 Empirical covariances & correlations: summary
  • The hypothesis according to which most of the volatility comes from fundamental information is hard to reconcile with the data. First, as discussed in section The role of impact fluctuations, this assumption does not enable one to rationalize why the square-root law, which applies to all metaorders (informed or not [25, 12, 46]) is proportional to volatility. Second, it predicts that the covariance of price returns and volume imbalances is proportional to $T$ independently of $a$, at odds with empirical data. As argued in [8, 49, 5], the most plausible hypothesis is that volatility stems from trading alone. In other words, the excess volatility puzzle has a microstructural origin.
@@ 8 Conclusion
In order to answer these questions, we have introduced a new theoretical framework to describe metaorders with different signs, sizes and durations, possibly correlated between themselves, which all impact prices as a square-root of volume (which we assume as an input) but with a subsequent time decay characterized by an exponent $\beta\neq\frac{1}{2}$, i.e. different from the one suggested by the classic propagator model [7, 5] or the LLOB model [17]. We proposed a generalized propagator model to account for such a feature.
@@
We then established that the power-law tailed distribution of metaorder durations is not sufficient to counteract impact decay that leads to price sub-diffusion. Rather, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders, which is indeed present in data. In fact, we conjecture that the intra- and cross-correlations between child orders decay roughly in the same manner, a feature that may be crucial for explaining the success of the construction of synthetic metaorders from public data [36]. Such a prediction is empirically falsifiable, and we look forward to more work in that direction using, e.g., the Tokyo Stock Exchange data of [45, 20].
@@
The existence of correlations between metaorders is therefore a crucial ingredient to recover price diffusion. The old debate between order splitting and “herding” that seemed to have been closed by several papers in favor of splitting [52, 45], is perhaps not so clear-cut. Such correlations could be due either to the fact that many participants use the same trading signals, or that copy-cat metaorders/unwinding market maker inventories correlate to past order flow, or else that some traders successfully predict the future behaviour of other participants. Note that within our story any predictive alpha signal manifests itself through autocorrelated metaorders (see also the discussion in [5], chapter 20).
@@
In view of the strongly fluctuating volumes $q$ of child orders, one quickly realizes that one needs to account for heterogeneity in the distribution of metaorder durations, and in the resulting decay of their impact, which we parametrized by two $q$-dependent exponents, $\mu_{q}$ and $\beta_{q}$, assumed to depend linearly on $\ell=\log q$. This feature allowed us to account semi-quantitatively for the way the moments of generalized volume imbalance scale with time $T$, and more importantly how the correlation between price changes and generalized volume imbalance scales with $T$. We predicted that the corresponding power-law should depend in a non-monotonic fashion on the parameter $a$ that allows one to put the same weight on all child orders ($a=0$) or overweight large orders ($a$ large), a behaviour clearly borne out by empirical data, see Fig. 7. We also predicted that the correlation between price changes and volume imbalances should display a maximum as a function of $a$ for fixed $T$, which again matches observations, see Fig. 9, with fitting parameters fixed with previously determined values. We found that stocks and futures appear to differ quite markedly in terms of these metrics, which could provide an interesting new way to characterize price formation mechanisms in different markets.
@@
Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We claim that our results strongly support the “Order-Driven” theory of markets, according to which it is the mechanical impact of trades, independently of any notion of “fundamental information”, that generates volatility in financial markets, a picture advocated in [8, 29, 5, 10] and, in a different context, in [24]. In particular, the Efficient Market Hypothesis, which posits that volatility is mostly due to the variation of “fundamental value”, cannot easily explain our results concerning the correlations between price changes and order flow imbalance.
@@
All these ideas are quantitatively confirmed using a numerical simulation of our model in a follow up of this work [maitrier2025subtle]. Such a numerical implementation will allow us in to (i) generate realistic synthetic datasets that capture the subtle interplay between order flows and prices, and (ii) validate on such synthetic data the method of metaorder reconstruction from the public tape recently proposed in [36].
Note added
After submission of this paper, we received the very interesting preprint of Muhle-Karbe et al., A unified theory of order flow, market impact, and volatility (January 30, 2026). The precise relation between our approach and theirs certainly deserves more investigations. We thank M. Rosenbaum for sharing his paper with us.